Hello,
I need help solving this problem, thanks.
Solve by elimination method.
2r-7s=11
7r+2s=65

Answer by stanbon(75887)

2r-7s=11
7r+2s=65
----------------
Multiply thru 1st by 7
Multiply thru 2nd by 2
----
14r - 49s = 77
14r + 4s = 130
-----------------
Subtract 1st from 2nd and solve for "s":
53s = 53
s = 1
----
Substitute and solve for "r":
2r -7*1 = 11
2r = 4
r = 2
-----
Cheers,
Stan H.